# On the wall-normal velocity variance in canonical wall-bounded turbulence

### v1

@@ 5 Summary
The amplitude and wavenumber of the spectrum peak scale with the local shear velocity $u_{\tau z}$ (Eq. 5) and dissipation-based length scale $\ell_{\epsilon}$ (Eq. 6). This scaling persists from the viscous sublayer to approximately half the boundary layer depth (Figs. 4 and 5). While these parameters are equivalent to the attached parameters $U_{\tau}$ and $z$ in the near-wall region at high Reynolds numbers (Fig. 3), the generalized definitions account also for trends at small $Re_{\tau}$ and across flow configurations. Further, $\ell_{\epsilon}$ transitions the spectrum directly into the inertial subrange which depends only on the local rate of dissipation (§3.2).
@@
The observed spectrum scaling results in the wall-normal variance depending on $u_{\tau z}$ and a local Reynolds number $Re_{\epsilon}=\ell_{\epsilon}/\eta$ (Fig. 6c). The dependence is approximated using a leading-order $Re_{\epsilon}^{-2/3}$ term derived from the inertial subrange and a higher-order empirical term to correct for the dissipative cutoff (Eq. 13). The resulting semi-emperical expression aligns closely with the present DNS and other variance measurements in the literature (Fig. 7).
@@
Extrapolating Eq. (11) to the limit $Re_{\tau}\to\infty$ suggests $B_{3}\approx 1.55\pm 0.1$, which is within the range of values previously reported and is closely aligned with empirical fits to atmospheric boundary layer measurements (Panofsky et al., 1977). However, the scaling with $u_{\tau z}$ rather than $U_{\tau}$ is new, and is attributed to the wall-normal motions being predominately “active” per the terminology of Townsend’s (1976) attached eddy hypothesis, which ties $w$ to the local Reynolds shear stress. The local scaling with $u_{\tau z}$ accounts for most of the difference in variance observed for ZPG and enclosed flows (Fig. 8), which have different stress profiles. While the constant stress profile in Eq. (2) is approximately achieved for ZPG flows in the high-$Re_{\tau}$ limit, the local scaling is required for pressure-driven flows to account for the linear total stress decay.
@@
The lowest wavenumbers in the $E_{ww}(k_{y})$ spectrum contribute relatively little to the total variance (Fig. 6a), but this contribution increases with wall-normal distance and is greater for ZPG flows (Figs. 4b and 5f). The motions at these wavenumbers are considered “inactive” owing to the absence of energy at the same wavenumbers in the shear stress spectrum (Fig. 9). The minor yet non-negligible dependence of the variance on these inactive motions, which changes with distance and flow configuration, leads to a range of possible $B_{3}$ values in the high-Reynolds-number limit. If so, the work here suggests there is not a single universal value for $B_{3}$ applicable to all wall-bounded flows.
@@ Acknowledgements.
All data are available from the original DNS studies (Sillero et al., 2013; Lee & Moser, 2015; Yao et al., 2023) and their corresponding datasets hosted online.

### v2

@@ 5 Summary
The amplitude and wavenumber of the spectrum peak are proportional to the local shear velocity $u_{\tau z}$ (Eq. 5) and dissipation-based length $\ell_{\epsilon}$ (Eq. 6). This dependence persists from the viscous sublayer to approximately half the boundary layer depth (Figs. 4 and 5). While these parameters are equivalent to the attached scales $U_{\tau}$ and $z$ in the near-wall region at high Reynolds numbers (Fig. 3), the generalized length $\ell_{\epsilon}$ accounts for trends at small $Re_{\tau}$, is applicable across the entire lower half of the boundary layer, and transitions the spectrum directly into the inertial subrange which depends only on the local rate of dissipation (§3.2).
@@
The observed trends in the spectra result in the wall-normal variance depending on $u_{\tau z}$ and a local Reynolds number $Re_{\epsilon}=\ell_{\epsilon}/\eta$ (Fig. 6c). The dependence is approximated using a leading-order $Re_{\epsilon}^{-2/3}$ term derived from the inertial subrange and a higher-order empirical term to correct for the dissipative cutoff (Eq. 13). The resulting semi-empirical expression aligns closely with the present DNS and other variance measurements in the literature (Fig. 7), noting that further investigation is required to verify and refine the empirical fit. For predicting the variance in practical applications, the dependencies $u_{\tau z}$ and $Re_{\epsilon}$ in Eq. (11) can be estimated as functions of $U_{\tau}$, $Re_{\tau}$, and $z/\delta$ that would differ between ZPG and pressure-driven flows.
@@
Extrapolating Eq. (11) to the limit $Re_{\tau}\to\infty$ suggests $B_{3}\approx 1.55\pm 0.1$, which is within the range of values previously reported and is closely aligned with empirical fits to atmospheric boundary layer measurements (Panofsky et al., 1977). The dependence of $B_{3}$ on $u_{\tau z}$ rather than $U_{\tau}$ alone is consistent with previous evidence (Tuerke and Jiménez, 2013; Lozano-Durán and Bae, 2019) and is attributed to the wall-normal motions being predominately ‘active’ per the terminology of Townsend’s (1976) attached eddy hypothesis, which ties $w$ to the local Reynolds shear stress. The hypothesis assumes a constant stress profile (Eq. 1.2), which is only achieved for ZPG flows in the high-$Re_{\tau}$ limit. The stress profile for each case, represented by $u_{\tau z}$, accounts for most of the differences in wall-normal variance observed across channel, pipe, ZPG-TBL, and plane Couette flows (Fig. 8). Achieving convincing scaling for second-order statistics has been an ongoing challenge owing to dependencies on both local and global flow properties that can differ across canonical cases (Marusic et al., 2010), and the present work shows that the wall-normal component of the variance is predominately governed by the local flow properties.
@@
The lowest wavenumbers in the $E_{ww}(k_{y})$ spectrum contribute relatively little to the total variance (Fig. 6a), but this contribution increases with wall-normal distance and is greater for ZPG flows (Figs. 4b and 5f). The motions at these wavenumbers are speculated to be ‘inactive’ owing to the absence of energy at the same wavenumbers in the shear stress spectrum (Fig. 9) and the fact that the low-wavenumber energy is poorly described by both distance-from-the-wall and local flow parameters (Fig. 4). The small yet non-negligible dependence of the variance on these inactive motions, which changes with distance and flow configuration, is a minor departure from Townsend’s hypothesis that considers wall-normal fluctuations to be purely ‘active’. While the present results support the hypotheses for attached eddies and active motions to first-order accuracy, these dissimilar low-wavenumber contributions suggest there is a small range of possible $B_{3}$ values in the high-Reynolds-number limit rather than a single universal value applicable to all canonical wall-bounded flows.
@@ Acknowledgements.
All data are available from the original DNS studies (Sillero et al., 2013; Lee and Moser, 2015; Yao et al., 2023) and their corresponding datasets hosted online.